EDITORIAL

Terry Goodman

We have received very positive comments concerning the changes we have made in the format of MJMS. Our Winter, 1996 edition was the first using the new format (size and cover design). We will continue to welcome your input concerning these changes. We are still considering the addition of a fourth issue each year and we encourage you to contact us with your opinion concerning this change.

The National Science Foundation is currently supporting several secondary mathematics curriculum development projects. These projects are designed to provide dramatically different mathematics curricular alternatives. Among these curriculum development projects are: CORE Plus (Western Michigan University), ARISE (CoMap), Interactive Mathematics Program (San Francisco State University), Math Connections (Connecticut Business and Industry Association), and SIMMS (University of Montana). Each project is seeking to develop a comprehensive program that is based on an integrated, problem–based model. Some common features of these projects are:

• a focus on applications that interest and motivate student investigation;

• inclusion of significant mathematical concepts, which are new to the secondary curriculum;

• course sequence based on an integrated model;

• requirement of teachers and students to assume roles of coach and co– investigator;

• use of alternative forms of evaluation

Teachers of undergraduate mathematics courses will be impacted by these curriculum reform efforts in several ways. First, high school graduates entering undergraduate mathematics programs will be coming from a secondary mathematics curriculum that is very different from the current, traditional curriculum. College/university mathematics faculty must consider how the undergraduate mathematics curriculum should be adjusted in light of alternative curricula at the secondary level. Secondly, those undergraduate mathematics faculty who help prepare preservice secondary mathematics teachers must consider changes needed in teacher preparation programs (content and methods) so that future teachers will be prepared to teach using alternative curriculum materials and instructional strategies.

There is currently much discussion concerning mathematics curriculum reform at the undergraduate level. Those who teach at this level should also begin to become familiar with curriculum reform at the secondary level so that reform efforts at both levels can be mutually supportive. In the future, MJMS will provide a forum in which discussion of mathematics curriculum reform and the related questions and issues can be discussed. We welcome your input. æ